OpenStudy (anonymous):

Find amount of money you will have after 5 years if 14,000 is invested in accounts of paying 8.25% interest compounded

OpenStudy (mathstudent55):

You haven't finished the problem. How often is the interest compounded?

OpenStudy (anonymous):

anually

OpenStudy (mathstudent55):

Ok. There is a formula for compound interest.

OpenStudy (anonymous):

do you want me to find it?

OpenStudy (mathstudent55):

\(A = P(1 + \dfrac{r}{n})^{nt} \) where A is future value P is the principal amount now r is the annual rate of interest written as a decimal n is the number of times the interest is compounded per year t is the number of years

OpenStudy (mathstudent55):

Just plug in all your values and evaluate A.

OpenStudy (anonymous):

is 14,000 the principal amount?

OpenStudy (funnyguy2):

Can not be answered unless you answer this: How often is the interest compounded?

OpenStudy (anonymous):

I'm not sure, how do I find out?

OpenStudy (anonymous):

??

OpenStudy (mathstudent55):

\(\Large A = 14,000(1 + \dfrac{0.0825}{1})^{1\times 5}\)

OpenStudy (mathstudent55):

\(\Large A = 14,000(1 + 0.0825)^{5}\) \(\Large A = 14,000(1.0825)^{5}\)

OpenStudy (anonymous):

The answer I got is 20809.78345

OpenStudy (mathstudent55):

Now raise 1.0825 to the 5th power an multiply by 14,000.

OpenStudy (mathstudent55):

Let me do it.

OpenStudy (mathstudent55):

Correct.